can you show step by step instructions on this problem? 30/x = 5/2
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

30/x = 5/2
----
Invert both sides to get:
x/30 = 2/5
Multiply both sides by 30 to solve for "x":
x = 30(2/5)
x = 6*2
x = 12
=================
Cheers,
Stan H.

Answer by Alan3354(69443)

30/x = 5/2
Cross multiply
30*2 = x*5
5x = 60
x = 12